OpenStudy (anonymous):

Simplify the expression.\[\sqrt{39}(\sqrt{6+ 7})\]

OpenStudy (anonymous):

First, we can simplify the 6+7 underneath the radical to get: \[\sqrt{39}\sqrt{13}\] Now, since both terms under the radicals are positive, we can multiply them together and put them under one radical: \[\sqrt{39\times13}\] or \[\sqrt{507}\] Try simplifying this further, if possible, and let me know if you'd like more help.

OpenStudy (anonymous):

from there wouldn't you do the square root of 507?!

OpenStudy (anonymous):

I'm sorry, the 6 should be outside the second radical sign, but in your question you asked it with the radical sign over the entire 7+6 expression, which means our answer is a little different. We need to distribute the rad 39 now to get: \[\sqrt{39\times6} + 7\sqrt{39} = \sqrt{234} + 7\sqrt{39}\]

OpenStudy (dan815):